OpenStudy (anonymous):

Which number is NOT contained in the domain of function f if f(x)=x+2/x+3 - 1/x??

OpenStudy (helder_edwin):

well. u have the function \[\large f(x)=\frac{x+2}{x+3}-\frac{1}{x}. \] u cannot have division by zero so u have to exclude the \(x\) such that \[\large x+3=0\qquad\text{and/or}\qquad x=0 \]

OpenStudy (anonymous):

i dont understand :(

OpenStudy (anonymous):

Another way of stating the question is, where is the function undefined?

OpenStudy (anonymous):

sorry, thats the way my teacher gave us the question. :/

OpenStudy (anonymous):

do you think you can help?

OpenStudy (anonymous):

I am just giving you an equivalent way of stating the question to help you understand it. Do you know how to find where the function is undefined?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know that division by zero is undefined?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the question is, for which values of \(x\) is there division by zero. It is for these values that the function is undefined. Now refer to helder_edwin's answer. He discusses this in more detail.